An electron is in two separate electric fields, one that is 1500 N/C West and one that is 2000N/C North. If the mass of an electron is 9.1 x 10^-31kg, find the acceleration of the electron.
1
Expert's answer
2019-06-24T09:15:49-0400

As long as both electric field vectors are perpendicular to each other, the resulting field can be obatined by means of a Pythagorean theorem:


"E=\sqrt{E_1^2 + E_2^2}"

One can also mention that direction of this resulting field is somewhere in the North-Western direction (not needed for the solution of the current problem).

The magnitude of the electrostatic force acting on the electron is


"F = eE"

which is oriented in the opposite to the field direction. As it is also the net force, one can apply the 2nd Newton's law:

"F = m a"

Hence, we obtain:


"a = \frac{e E}{m} = \frac{ e \sqrt{E_1^2 + E_2^2}}{m}"

Substituting the numerical values, we obtain:


"a = \frac{1.6 \cdot 10^{-19} \sqrt{1500^2 + 2000^2}}{9.1 \cdot 10^{-31}} \approx 4.4 \cdot 10^{14} \, m/s^2"

Answer: 4.4*1014 m/s2